OpenStudy (kkutie7):

A roller coaster moves 200 ft. horizontally, hen rises 135 ft at an angle 30degrees above the horizontal. It then travels 135 ft at an angle of 40 degrees below horizontal. What is its displacement from its starting point at the end of this movement? Use graphical techniques and also find the magnitude and direction.

OpenStudy (kkutie7):

Answer: 537.26 ft horizontal and 48.22 ft vertical, 539.41 ft at 5.13◦ how do I get here?

OpenStudy (anonymous):

Draw a picture: http://i279.photobucket.com/albums/kk154... x_tot = 200 + [135cos(30°)] + [135cos(-40°)] x_tot = 420.3294293 ft y_tot = [135sin(30°)] + [135sin(-40°)] y_tot = -19.27632731 ft Displacement = √(x² + y²) Displacement = √[(420.3294293)² + (-19.27632731)²] Displacement = 420.771204 ft Displacement = 421 ft tanθ = (y / x) tanθ = (-19.27632731 / 420.3294293) θ = arctan (-0.0458600468) θ = -2.625747386° θ = -2.63° Displacement = 421 ft at 2.63° below the horizontal.

OpenStudy (anonymous):

Displacement = 421 ft at 2.63° below the horizontal.

OpenStudy (anonymous):



OpenStudy (kkutie7):

Link is broken =(

OpenStudy (anonymous):

:( sorry

OpenStudy (kkutie7):

Does it look something like this?